# The radius of statistical efficiency

### v1

@@ Abstract
Classical results in asymptotic statistics show that the Fisher information matrix controls the difficulty of estimating a statistical model from observed data. In this work, we introduce a companion measure of robustness of an estimation problem: the radius of statistical efficiency (RSE) is the size of the smallest perturbation to the problem data that renders the Fisher information matrix singular. We compute RSE up to numerical constants for a variety of test bed problems, including principal component analysis, generalized linear models, phase retrieval, bilinear sensing, and matrix completion. In all cases, the RSE quantifies the compatibility between the covariance of the population data and the latent model parameter. Interestingly, we observe a precise reciprocal relationship between RSE and the intrinsic complexity/sensitivity of the problem instance, paralleling the classical Eckart–Young theorem in numerical analysis.
@@ 1 Introduction
A central theme in computational mathematics is that the numerical difficulty of solving a given problem is closely linked to both $(i)$ the sensitivity of its solution to perturbations and $(ii)$ the shortest distance of the problem to an ill-posed instance. As a rudimentary example, consider solving an $m\times d$ linear system ${A}{x}={b}$. The celebrated Eckart–Young theorem asserts the equality:
$\underbrace{\min_{B\in\mathbf{R}^{m\times d}}\left\{\|A-B\|_{F}\mid B~{}\rm{is~{}singular}\right\}}_{\rm distance~{}to~{}ill-posedness}\quad=\underbrace{\sigma_{\min}({A})}_{\rm difficulty/sensitivity}.$ (1)
Although the proof is elementary, the conclusion is intriguing since it equates two conceptually distinct quantities. Namely, the reciprocal of the minimal singular value $1/\sigma_{\min}({A})$ is classically known to control both the numerical difficulty of solving the linear system ${A}{x}={b}$ and the Lipschitz stability of the solution to perturbations in ${b}$. In contrast, the left side of the equation (1) is geometric; it measures the smallest perturbation to the data that renders the problem ill-posed.
@@
We will see that in a certain precise sense the answer is indeed yes for a wide class of problems. The starting point for our development is that the statistical difficulty of estimation is tightly controlled by (quantities akin to) the Fisher information matrix for maximum likelihood estimation. This connection is made precise for example by the Cramér-Rao lower bound [19, 53] and the local asymptotic minimax theory of Hájek and Le Cam [39, 65, 35]. From an optimization viewpoint, the minimal eigenvalue of the Fisher information matrix is closely related to the quadratic growth constant of the objective function modeling the learning problem at hand. In particular, (near-) singularity of this matrix signifies that the problem is ill-conditioned. Inspired by this observation, we introduce a new measure of robustness associated to an estimation task: the radius of statistical efficiency (RSE) is the size of the smallest perturbation to the problem data that renders the Fisher information matrix singular. Thus large RSE signifies existence of a large neighborhood around the problem instance comprised only of well-posed problems.
@@
where ${x}\in\mathbf{R}^{d}$ is drawn from a probability distribution $\mathsf{D}$ and $\varepsilon$ is zero-mean noise vector that is independent of ${x}$. The standard approach to this task is to minimize the mean-squared error
$\min_{{\beta}\in{\mathbf{R}}^{d}}~{}\tfrac{1}{2}\operatorname{\mathbb{E}}_{{x},y}\left(\langle{x},{\beta}\rangle-y\right)^{2}.$ (2)
Classical results show that the asymptotic performance of estimators for this problem is tightly controlled by ${\Sigma}^{-1}$ where ${\Sigma}:=\operatorname{\mathbb{E}}_{{x}\sim\mathsf{D}}{x}{x}^{\top}$ is the second moment matrix of the population. The closer the matrix $\Sigma$ is to being singular, the more challenging the problem (2) is to solve, requiring a higher number of samples. Seeking to estimate a neighborhood of well-posed problems around $\mathsf{D}$, the RSE is defined to be the minimal Wasserstein-2 distance from $\mathsf{D}$ to distributions with a singular second moment matrix. We will see that for linear regression (2), RSE is simple to compute:
$\operatorname{RSE}(\mathsf{D})=\sqrt{\lambda_{\min}\left({\Sigma}\right)}.$ (3)
That is, the simplest measure of ill-conditioning of the target problem $1/\sqrt{\lambda_{\min}\left({\Sigma}\right)}$ has a geometric interpretation as the reciprocal of the distance to the nearest ill-posed problem. The representation of RSE in (3) is not specific to linear regression and holds much more generally for (quasi) maximum likelihood estimation [46] with strongly convex cumulant functions.
@@
$\max_{{R}\in{\rm Gr}(q,d)}~{}\mathop{\operatorname{\mathbb{E}}}_{{x}\sim\mathsf{D}}\|{R}{x}\|^{2}_{2},$ (4)
where the Grassmannian manifold ${\rm Gr}(q,d)$ consists of all orthogonal projection matrices ${R}\in\mathbf{R}^{d\times d}$ onto $q$-dimensional subspaces of $\mathbf{R}^{d}$. The column space of the optimal matrix $R$ is called the top $q$ principal subspace. Intuitively, the hardness of the problem is governed by the gap $\lambda_{q}-\lambda_{q+1}$ between the $q$’th and $(q+1)$’th eigenvalues of $\Sigma=\operatorname{\mathbb{E}}_{\mathsf{D}}{{x}{x}^{\top}}$: the smaller the gap is, the higher the number of samples required for estimation, and this can indeed be made rigorous. Again, RSE by definition is the minimal Wasserstein-2 distance from $\mathsf{D}$ to a distribution with covariance having equal $q$’th and $(q+1)$’st eigenvalues. We will show that RSE admits the simple form
$\operatorname{RSE}(\mathsf{D})=\tfrac{1}{\sqrt{2}}\left(\sqrt{\lambda_{q}({\Sigma})}-\sqrt{\lambda_{q+1}({\Sigma})}\right)\!.$ (5)
@@
$\min_{{\beta}\in\mathcal{M}}~{}f({\beta})\quad\text{where}\quad f({\beta})=\mathop{\operatorname{\mathbb{E}}}_{{z}\sim\mathsf{D}}\ell({\beta};{z}).$ $\mathtt{SP}(\mathsf{D})$
@@
Here ${z}$ is data drawn from a distribution $\mathsf{D}$, the function $\ell({\beta};\cdot)$ is a loss parameterized by ${\beta}$, and $\mathcal{M}\subset\mathbf{R}^{d}$ is a smooth manifold of allowable model parameters. For example, $\mathtt{SP}(\mathsf{D})$ may correspond to least-squares regression or maximum likelihood estimation. In both cases, the asymptotic efficiency of estimators for finding the minimizer ${\beta}_{\star}$ of $\mathtt{SP}(\mathsf{D})$ is tightly controlled by the following matrix akin to Fisher information:
$\mathcal{I}({\beta}_{\star},\mathsf{D})=P_{\mathcal{T}}\nabla_{{\beta}{\beta}}^{2}\mathcal{L}({\beta}_{\star},{z},{\lambda}_{\star})P_{\mathcal{T}}.$ (6)
@@
Here, $P_{\mathcal{T}}$ is the projection onto the tangent space of $\mathcal{M}$ at ${\beta}_{\star}$ and $\mathcal{L}({\beta}_{\star},{z},{\lambda}_{\star})$ is the Lagrangian function for $\mathtt{SP}(\mathsf{D})$ with optimal multiplier ${\lambda}_{\star}$. For simplicity, we will abuse notation and call $\mathcal{I}({\beta}_{\star},\mathsf{D})$ the Fisher information matrix. The matrix $\mathcal{I}({\beta}_{\star},\mathsf{D})$ plays a central role in estimation, as highlighted by the lower bounds of Cramér-Rao and Hájek-Le Cam [39, 65], as well as their recent extensions to stochastic optimization of Duchi-Ruan [35]. Moreover, the minimal nonzero eigenvalue of $\mathcal{I}({\beta}_{\star},\mathsf{D})$ controls both the coefficient of quadratic growth of the objective function in $\mathtt{SP}(\mathsf{D})$ and the Lipschitz stability of the solution under linear perturbations. In particular, the problem $\mathtt{SP}(\mathsf{D})$ becomes ill-conditioned when the minimal eigenvalue of $\mathcal{I}({\beta}_{\star},\mathsf{D})$ is small.
@@
In summary, the matrix $\mathcal{I}({\beta}_{\star},\mathsf{D})$ tightly controls the difficulty of solving $\mathtt{SP}(\mathsf{D})$. Consequently, it is appealing to consider as a measure of robustness of $\mathtt{SP}(\mathsf{D})$ the size of the smallest perturbation to the data $\mathsf{D}$, say in the Wasserstein-2 distance $W_{2}(\cdot,\cdot)$, that renders the Fisher information matrix singular. This is the viewpoint we explore in the current work, and call this quantity the radius of statistical efficiency (RSE). We choose to use the Wasserstein-2 distance in the definition of RSE, as opposed to other metrics on measures, because it leads to concise and easily interpretable estimates in examples. In the rest of the paper, we study basic properties of RSE and compute it up to numerical constants for a variety of test bed problems: generalized linear models, PCA, phase retrieval, blind deconvolution, and matrix completion. In all cases, the RSE translates intuitive measures of “well-posedness” into quantified neighborhoods of stable problems. Moreover, in all cases we show a reciprocal relationship between the minimal eigenvalue of $\mathcal{I}({\beta}_{\star},\mathsf{D})$ and RSE, thereby paralleling the Eckart–Young theorem in numerical analysis and optimization.
@@ 1.1 Related work
There is a rich literature on minimax lower bounds in statistical estimation problems; we refer the reader to [68, Chapter 15] for a detailed treatment. Typical results of this type lower-bound the performance of any statistical procedure on a worst-case instance for that procedure. Minimax lower bounds can be quite loose as they do not consider the complexity of the particular problem that one is trying to solve but rather that of an entire problem class to which it belongs. More precise local minimax lower bounds, as developed by Hájek and Le Cam [39, 65], provide much finer problem-specific guarantees. Simply put, a single object akin to the Fisher information matrix controls both the difficulty of estimation from finitely many samples and the stability of the model parameters to perturbation of the density. Extensions of this theory to stochastic nonlinear programming were developed by Duchi and Ruan [35] and extended to decision-dependent problems in [21] and to a wider class of (partly smooth) problems in [26]. In particular, it is known that popular algorithms such as sample average approximation [65] and stochastic gradient descent with iterate averaging [51, 26] match asymptotic local lower-bound, and are therefore asymptotically optimal. Weaker ad hoc results, based on the Cramér-Rao lower bound, have been established for a handful of problems [48, 44, 62, 3, 4].
@@
Classical numerical analysis literature emphasizes the close interplay between efficiency of numerical algorithms and their sensitivity to perturbation. Namely, problems with solutions that change rapidly due to small perturbation are typically difficult to solve. Examples of this phenomenon abound in computational mathematics; e.g. eigenvalue problems and polynomial equations [29, 13] and optimization [31, 32, 57]. Motivated by this observation, Demmel in [28] introduced a new robustness measure, called the radius of regularity, which measures the size of a neighborhood around a problem instance within which all other problem instances are stable. A larger neighborhood thereby signifies a more robust problem instance. Estimates on the radius of regularity have now been computed for a wealth of computational problems; e.g. solving polynomial systems [12, 13], linear and conic programming [56, 55, 36, 49], and nonlinear optimization [31]. The radius of statistical efficiency, introduced here, serves as a direct analogue for statistical estimation.
@@
Several condition numbers—controlling the convergence of first-order methods—are closely related to notions of strong identifiability, e.g., the restricted isometry property (RIP), in the context of statistical recovery problems [8, 16, 45, 14, 15, 30]. A few works [58, 71, 5] have established connections between these notions of strong identifiability and a suitably-defined radius to ill-posed instances. In particular, [58] established a connection between Renegar’s conic distance to infeasibility [56] and the null space property [18] in compressed sensing. In a similar vein, [71] linked the $\ell_{1}$-distance to ill-posed problems and the RIP for generalized rank-one matrix completion. Finally, [5] defined a condition number for the LASSO variable selection problem via the reciprocal of the distance to ill-posedness, designed an algorithm whose complexity depends solely on this condition number, and proved an impossibility result for instances with infinite condition number. The radius of statistical efficiency, defined in this work, is distinct from and complementary to these metrics.
@@
The basic question we explore is the relationship between the minimal eigenvalue of the Fisher information matrix (complexity) and the distance to the set where this eigenvalue is zero (RSE). The theory of error bounds exactly addresses questions of this type; namely when does the function’s value polynomially bound the distance to the set of minimizers. See for example the authoritative monographs on the subject, [20, Chapter 8] and [38, Chapter 3]. Indeed, Demmel’s original work [28] makes heavy use of this interpretation. We explore this path here as well when developing infinitesimal characterizations of RSE in Theorem 2. The added complication is that the functions we deal with are defined over a complete metric space, and therefore the techniques we use rely on variational principles (á la Ekeland) and computations of the slope.

### v2

@@ Abstract
Classical results in asymptotic statistics show that the Fisher information matrix controls the difficulty of estimating a statistical model from observed data. In this work, we introduce a companion measure of robustness of an estimation problem: the radius of statistical efficiency (RSE) is the size of the smallest perturbation to the problem data that renders the Fisher information matrix singular. We compute RSE up to numerical constants for a variety of testbed problems, including principal component analysis, generalized linear models, phase retrieval, bilinear sensing, and matrix completion. Interestingly, we observe a precise reciprocal relationship between RSE and the intrinsic complexity/sensitivity of the problem instance, paralleling the classical Eckart–Young theorem in numerical analysis. To establish our results, we develop theory for spectral functions of measures that extends well-known results from matrix analysis and eigenvalue optimization—a contribution that may be of interest beyond our immediate findings.
Keywords— radius theorem, conditioning, statistical complexity, Eckart–Young theorem, stability, robustness, Fisher information matrix, local minimax lower bounds.
MSC numbers— 90C15, 49K40, 62F12, 90C31.
@@ 1 Introduction
A central theme in computational mathematics is that the numerical difficulty of solving a given problem is closely linked to both $(i)$ the sensitivity of its solution to perturbations and $(ii)$ the shortest distance of the problem to an ill-posed instance. As a rudimentary example, consider solving an $m\times d$ linear system Ax=b{A}{x}={b}. The celebrated Eckart–Young theorem asserts the equality
$\underbrace{\min_{B\in\mathbf{R}^{m\times d}}\left\{\|A-B\|_{F}\mid B~\rm{is~singular}\right\}}_{\clap{\scriptsize{\rm distancetoill-posedness}}}\kern 5.0pt=\qquad\underbrace{\sigma_{\min}({A})}_{\clap{\scriptsize{\rm difficulty/sensitivity}}}.$ (1)
Although the proof is elementary, the conclusion is intriguing since it equates two conceptually distinct quantities. Namely, the reciprocal of the minimal singular value $1/\sigma_{\min}({A})$ is classically known to control both the numerical difficulty of solving the linear system Ax=b{A}{x}={b} and the Lipschitz stability of the solution to perturbations in b{b}. In contrast, the left side of the equation (1) is geometric; it measures the smallest perturbation to the data that renders the problem ill-posed.
@@
As stated, this question is somewhat ill-posed. Unlike in the linear system case where there is a clear parameterization of the problem, statistical inference problems take many different forms with varying structure. Instead of pursuing overgeneralized answers, we demonstrate that for a specific class of problems resembling maximum likelihood estimation, a precise relationship can indeed be established. The starting point for our development is that the statistical difficulty of estimation is tightly controlled by (quantities akin to) the Fisher information matrix for maximum likelihood estimation. This connection is made precise for example by the Cramér-Rao lower bound [19, 56] and the local asymptotic minimax theory of Hájek and Le Cam [40, 69, 35]. From an optimization viewpoint, the minimal eigenvalue of the Fisher information matrix is closely related to the quadratic growth constant of the objective function modeling the learning problem at hand. In particular, (near-)singularity of this matrix signifies that the problem is ill-conditioned. Inspired by this observation, we introduce a new measure of robustness associated to an estimation task: the radius of statistical efficiency (RSE) is the size of the smallest perturbation to the problem data that renders the Fisher information matrix singular. Thus large RSE signifies existence of a large neighborhood around the problem instance comprised only of well-posed problems.
@@
where x{x} is a random vector in $\mathbf{R}^{d}$ with probability distribution $\mathsf{D}_{x}$ and $\varepsilon$ is zero-mean noise that is independent of x{x}. The standard approach to this task is to minimize the mean squared error
$\min_{{\beta}\in{\mathbf{R}}^{d}}\,\mathop{\operatorname{\mathbb{E}}}_{(x,y)\sim\mathsf{D}}\tfrac{1}{2}\big(\langle{x},{\beta}\rangle-y\big)^{2}.$ (2)
Classical results show that the asymptotic performance of estimators for this problem is tightly controlled by ${\Sigma}^{-1}$, where ${\Sigma}:=\operatorname{\mathbb{E}}_{\mathsf{D}_{x}}[{x}{x}^{\top}]$ is the second-moment matrix of the population. The closer the matrix $\Sigma$ is to being singular, the more challenging the problem (2) is to solve, requiring a higher number of samples. Seeking to estimate a neighborhood of well-posed problems around $\mathsf{D}$, the RSE is defined to be the minimal Wasserstein-2 distance from $\mathsf{D}_{x}$ to distributions with a singular second-moment matrix. We will see that for linear regression (2), RSE is simple to compute:
$\operatorname{RSE}(\mathsf{D})=\sqrt{\lambda_{\min}({\Sigma})}.$ (3)
That is, the simplest measure of ill-conditioning of the target problem $1/\lambda_{\min}({\Sigma})$ has a geometric interpretation as the reciprocal of the squared distance to the nearest ill-posed problem. The representation of RSE in (3) is not specific to linear regression and holds much more generally for (quasi) maximum likelihood estimation [48] with strongly convex cumulant functions.
@@
$\max_{{R}\in{\rm Gr}(q,d)}\,\mathop{\operatorname{\mathbb{E}}}_{{x}\sim\mathsf{D}}\|{R}{x}\|^{2}_{2},$ (4)
where the Grassmannian manifold ${\rm Gr}(q,d)$ consists of all orthogonal projection matrices ${R}\in\mathbf{R}^{d\times d}$ onto $q$-dimensional subspaces of $\mathbf{R}^{d}$. The column space of an optimal matrix $R$ is called the top $q$ principal subspace. Intuitively, the hardness of the problem is governed by the gap between the $q$’th and $(q+1)$’st eigenvalues of $\Sigma:={\operatorname{\mathbb{E}}}_{\mathsf{D}}{[xx^{\top}]}$—the smaller the gap, the higher the number of samples required for estimation, and this can indeed be made rigorous. We define the RSE for PCA to be the minimal Wasserstein-2 distance from $\mathsf{D}$ to a centered distribution with covariance having equal $q$’th and $(q+1)$’st eigenvalues. In this case, RSE admits the simple form
$\operatorname{RSE}(\mathsf{D})=\tfrac{1}{\sqrt{2}}\bigg(\sqrt{\lambda_{q}({\Sigma})}-\sqrt{\lambda_{q+1}({\Sigma})}\bigg).$ (5)
@@
$\min_{{\beta}\in\mathcal{M}}\,f({\beta})\quad\text{where}\quad f({\beta})=\mathop{\operatorname{\mathbb{E}}}_{{z}\sim\mathsf{D}}\ell({\beta};{z}).$ $\mathtt{SP}(\mathsf{D})$
@@
Here z{z} is data drawn from a distribution $\mathsf{D}$, the function $\ell({\beta};\cdot)$ is a loss parameterized by $\beta$}, and $\mathcal{M}\subset\mathbf{R}^{d}$ is a $C^{2}$ embedded submanifold comprised of allowable model parameters. For example, $\mathtt{SP}(\mathsf{D})$ may correspond to least-squares regression or maximum likelihood estimation. In both cases, the asymptotic efficiency of estimators for finding the minimizer $\beta$$_{\star}$ of $\mathtt{SP}(\mathsf{D})$ is tightly controlled by the following matrix akin to Fisher information:
$\mathcal{I}({\beta}_{\star},\mathsf{D})=\mathsf{P}_{\mathcal{T}}\nabla_{{\beta}{\beta}}^{2}\mathcal{L}({\beta}_{\star},{\lambda}_{\star})\mathsf{P}_{\mathcal{T}}.$ (6)
@@
Here $\mathsf{P}_{\mathcal{T}}$ is the projection onto the tangent space $\mathcal{T}$} of $\mathcal{M}$ at $\beta$$_{\star}$ and $\mathcal{L}$ is the Lagrangian function for $\mathtt{SP}(\mathsf{D})$ with optimal multipliers $\lambda$$_{\star}$. For simplicity, we will abuse notation and call $\mathcal{I}({\beta}_{\star},\mathsf{D})$ the Fisher information matrix. The matrix $\mathcal{I}({\beta}_{\star},\mathsf{D})$ plays a central role in estimation, as highlighted by the lower bounds of Cramér-Rao and Hájek-Le Cam [40, 69], as well as their recent extensions to stochastic optimization of Duchi-Ruan [35]. [^1] In general, these lower bounds consider a slightly more complicated matrix (see (8) in Section 2), but for the examples considered in this paper it reduces to (a scalar multiple of) the pseudoinverse of the matrix $\mathcal{I}({\beta}_{\star},\mathsf{D})$. Moreover, the minimal nonzero eigenvalue of $\mathcal{I}({\beta}_{\star},\mathsf{D})$ controls both the coefficient of quadratic growth of the objective function in $\mathtt{SP}(\mathsf{D})$ and the Lipschitz stability of the solution under linear perturbations. In particular, the problem $\mathtt{SP}(\mathsf{D})$ becomes ill-conditioned when the minimal eigenvalue of $\mathcal{I}({\beta}_{\star},\mathsf{D})|_{\mathcal{T}}$ is small.
@@
In summary, the matrix $\mathcal{I}({\beta}_{\star},\mathsf{D})$ tightly controls the difficulty of solving $\mathtt{SP}(\mathsf{D})$. Consequently, it is appealing to consider as a measure of robustness of $\mathtt{SP}(\mathsf{D})$ the size of the smallest perturbation to the data $\mathsf{D}$, say in the Wasserstein-2 distance $W_{2}(\cdot,\cdot)$, that renders the Fisher information matrix singular. This is the viewpoint we explore in the current work, and we call this quantity the radius of statistical efficiency (RSE). We choose to use the Wasserstein-2 distance in the definition of RSE, as opposed to other metrics on measures, because it leads to concise and easily interpretable estimates in examples. Indeed, the core machinery underlying our arguments can be summarized via the following observation. For most problems we study, the set of distributions with singular Fisher information matrix takes the form $\mathcal{P}_{\mathcal{C}}=\{\mu\mid\Sigma_{\mu}\in\mathcal{C}\}$, where $\Sigma_{\mu}$ denotes the second-moment matrix of the distribution $\mu$ and $\mathcal{C}$ is some given set. For each example, we establish a reduction akin to
$W_{2}(\mathsf{D},\mathcal{P}_{\mathcal{C}})=W_{2}(\Sigma_{\mathsf{D}},\mathcal{C}),$
where $W_{2}$ on the right-hand side denotes the Bures-Wasserstein distance between positive semidefinite matrices. Thus, computing the Wasserstein-2 distance reduces to an easier finite-dimensional problem; we defer a detailed discussion to Appendix A. In the rest of the paper, we leverage this strategy to study basic properties of RSE and compute it up to numerical constants for a variety of testbed problems: generalized linear models, PCA, phase retrieval, blind deconvolution, and matrix completion. In all cases, the RSE translates intuitive measures of “well-posedness” into quantified neighborhoods of stable problems. Moreover, in all cases we show a reciprocal relationship between the minimal eigenvalue of $\mathcal{I}({\beta}_{\star},\mathsf{D})$ and RSE, thereby paralleling the Eckart–Young theorem in numerical analysis and optimization.
@@ 1.1 Related work
There is a rich literature on minimax lower bounds in statistical estimation problems; we refer the reader to [72, Chapter 15] for a detailed treatment. Typical results of this type lower-bound the performance of any statistical procedure on a worst-case instance for that procedure. Minimax lower bounds can be quite loose as they do not consider the complexity of the particular problem that one is trying to solve but rather that of an entire problem class to which it belongs. More precise local minimax lower bounds, as developed by Hájek and Le Cam [40, 69], provide much finer problem-specific guarantees. Simply put, a single object akin to the Fisher information matrix controls both the difficulty of estimation from finitely many samples and the stability of the model parameters to perturbation of the density. Extensions of this theory to stochastic nonlinear programming were developed by Duchi and Ruan [35] and extended to decision-dependent problems in [21] and to a wider class of (partly smooth) problems in [26]. In particular, it is known that popular algorithms such as sample average approximation [69] and stochastic gradient descent with iterate averaging [54, 26] match asymptotic local minimax lower bounds, and are therefore asymptotically optimal. Weaker ad hoc results, based on the Cramér-Rao lower bound, have been established for a handful of problems [50, 46, 66, 3, 4]. Our results are different in nature—we do not derive local minimax lower bounds. Instead, we establish a relationship between the quantities appearing in these lower bounds and the distance to a set of ill-conditioned instances.
@@
Classical numerical analysis literature emphasizes the close interplay between efficiency of numerical algorithms and their sensitivity to perturbation. Namely, problems with solutions that change rapidly due to small perturbation are typically difficult to solve. Examples of this phenomenon abound in computational mathematics; e.g., eigenvalue problems and polynomial equations [28, 13] and optimization [32, 31, 60]. Motivated by this observation, Demmel in [29] introduced a new robustness measure, called the radius of regularity, which measures the size of a neighborhood around a problem instance within which all other problem instances are stable. A larger neighborhood thereby signifies a more robust problem instance. Estimates on the radius of regularity have now been computed for a wealth of computational problems; e.g., solving polynomial systems [12, 13], linear and conic programming [59, 58, 36, 52], and nonlinear optimization [32]. The radius of statistical efficiency, introduced here, serves as a direct analogue for statistical estimation.
@@
Several condition numbers—controlling the convergence of first-order methods—are closely related to notions of strong identifiability, e.g., the restricted isometry property (RIP), in the context of statistical recovery problems [8, 16, 47, 14, 15, 30]. A few works [61, 75, 5] have established connections between these notions of strong identifiability and a suitably-defined radius to ill-posed instances. In particular, [61] established a connection between Renegar’s conic distance to infeasibility [59] and the nullspace property [18] in compressed sensing. In a similar vein, [75] linked the $\ell_{1}$ distance to ill-posed problems and the RIP for generalized rank-one matrix completion. Finally, [5] defined a condition number for the LASSO variable selection problem via the reciprocal of the distance to ill-posedness, designed an algorithm whose complexity depends solely on this condition number, and proved an impossibility result for instances with infinite condition number. The radius of statistical efficiency, defined in this work, is distinct from and complementary to these metrics.
@@
The basic question we explore is the relationship between the minimal eigenvalue of the Fisher information matrix (complexity) and the distance to the set where this eigenvalue is zero (RSE). The theory of error bounds exactly addresses questions of this type: namely, when does a function’s value polynomially bound the distance to its set of minimizers? See, for example, the authoritative monographs on the subject [20, Chapter 8] and [39, Chapter 3]. Indeed, Demmel’s original work [29] makes heavy use of this interpretation. We explore this path here as well when developing infinitesimal characterizations of RSE in Theorem 2. The added complication in our development is that we must deal with functions defined over a metric space rather than a Euclidean space, and therefore the techniques we use rely on variational principles (á la Ekeland) and computations of the slope.